X, Y and Z are all gases that behave ideally and react according to the equation shown.
X(g) + 2Y(g) → 2Z(g)
When 3.0 mol of X and 3.0 mol of Y are placed inside a container with a volume of 1.0 dm3
, they
react to form the maximum amount of Z.
The final temperature of the reaction vessel is 120 °C.
What is the final pressure inside the reaction vessel?
1
Expert's answer
2020-11-02T08:50:33-0500

X(g) + 2Y(g) → 2Z(g)

When 3 mol Y reacts it forms 3 mol Z and 3/2 = 1.5 mol of X left

So, totally we have:

3 mol Z + 1.5 mol X = 4.5 mol of gases

PV = nRT

Where

n = 4.5 mol

V = 1 dm3 (L) = 0.001 m3

R = 8.314 J/mol K

T = 120 + 273.15 C = 393.15 K

Final pressure:

P = nRT / V

P = (4.5 x 8.314 x 393.15) / 0.001 = 14708920 Pa = 14.7 MPa